# Mirror Descent Algorithms with Nearly Dimension-Independent Rates for Differentially-Private Stochastic Saddle-Point Problems

### v1

@@ Appendix F Missing details from Section 5
$\quad+2\tau q\|\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t})\|_{\infty}^{2}+2\tau\sum_{k=t_{0}}^{t-1}\mathbb{E}\left[\|\nabla f(\hat{w}^{t_{0}};B^{k})-\nabla f(w^{t_{0}};B^{k})\|_{\infty}^{2}\right]$
$\lesssim\begin{cases}\frac{L_{1}}{\sqrt{K}}+q\tau\big{[}\frac{L_{1}^{2}}{\sqrt{\log(d_{x})}K^{3/2}}+\frac{(L_{0}^{2}+L_{1}^{2})\sqrt{\log(d_{x})}}{\sqrt{K}}\big{]}&\text{without second order smoothnes}\\ \frac{L_{2}}{K}+q\tau\big{[}\frac{L_{2}^{2}}{K^{2}}+\frac{L_{1}^{2}\log(d_{x})}{K}\big{]}&\text{with second order smoothness}\end{cases}$
And recall that for $t\leq q$
$\mathbb{E}[\langle\nabla F_{\mathcal{D}}(w^{t})-\nabla f(\hat{w}^{t};B^{t}),x^{t}-x\rangle]\leq\begin{cases}\frac{8L_{1}}{\sqrt{K}}&\text{without second order smoothnes}\\ \frac{4L_{2}}{K}&\text{with second order smoothness}\end{cases}.$
Finally, we plug this bounds into our excess risk bound. We start considering the case of second order smooth functions:
$\mathbb{E}[F_{\mathcal{D}}(w^{T})-F_{\mathcal{D}}(x)]$
$\leq\frac{\log(d_{x})}{\tau T}+\frac{\tau L_{0}^{2}}{2}+\frac{1}{T}\sum_{t\in[T]}\mathbb{E}\left[\langle\nabla F_{\mathcal{D}}(w^{t})-g^{t},x^{t}-x\rangle\right]$
$\lesssim\frac{\log(d_{x})}{\tau T}+\tau L_{0}^{2}+\frac{1}{T}\sum_{t\in[T]}\left(\frac{L_{2}}{K}+\mathbbm{1}_{(t>q)}\left[\frac{4L_{1}q}{q\lfloor t/q\rfloor+1}+q\frac{L_{2}^{2}\tau}{K^{2}}+q\frac{L_{1}^{2}\log(d_{x})\tau}{K}\right]\right)$
$\lesssim\frac{\log(d_{x})}{\tau T}+\tau\Big{[}L_{0}^{2}+\frac{L_{1}^{2}q\log(d_{x})}{K}+\frac{L_{2}^{2}q}{K^{2}}\big{]}+\frac{L_{2}}{K}+\frac{1}{T}\sum_{t=q+1}^{T}\left[\frac{L_{1}q}{q\lfloor t/q\rfloor+1}\right]$
$\lesssim\frac{\log(d_{x})}{\tau T}+\tau\Big{[}L_{0}^{2}+\frac{L_{1}^{2}q\log(d_{x})}{K}+\frac{L_{2}^{2}q}{K^{2}}\big{]}+\frac{L_{2}}{K}+\frac{L_{1}q\log(T/q)}{T},$ (5)
where the last line follows from $\sum_{t=q+1}^{T}\left[\frac{1}{q\lfloor t/q\rfloor+1}\right]\leq q\left(\frac{1}{q}+\frac{1}{2q}+...+\frac{1}{(T/q)q}\right)\lesssim\log\Big{(}\frac{T}{q}\Big{)}$. Now, by Theorem F.1, and recalling that the batch size is $B=n/T$, we have a constraint on $\tau>0$ in order to satisfy privacy. We also have the constraint of $\psi\leq 1$, which holds true only when $\tau\leq 1/(8qL_{0})$. Optimizing (5) under these constraints, we have that we should choose
$\tau\eqsim\min\left\{\left(\sqrt{\frac{\log(d_{x})}{(L_{0}^{2}+L_{1}^{2}q\log(d_{x})/K+L_{2}^{2}q/K^{2})T}}\right),\frac{1}{L_{0}q},\frac{n\varepsilon}{TL_{0}\sqrt{(TK/q+qK)\log(1/\delta)}}\right\},$
we obtain
@@
$\mathbb{E}[F_{\mathcal{D}}(\hat{w}^{T})-F_{\mathcal{D}}(x)]$ $\lesssim(L_{0}+L_{1})\sqrt{\frac{\log(d_{x})}{T}}+\frac{L_{2}}{T^{3/4}\log(d_{x})^{1/4}}+\frac{L_{1}+L_{2}}{\sqrt{T\log(d_{x})}}$
$+\frac{L_{1}\log(\sqrt{n\log(d_{x})})}{\sqrt{T\log(d_{x})}}+\frac{L_{0}\sqrt{T\log(1/\delta)}\log^{3/2}(d_{x})}{n\varepsilon}.$
$\lesssim(L_{0}+L_{1})\sqrt{\frac{\log(d_{x})}{T}}+\frac{L_{2}}{\sqrt{T}\log(d_{x})^{1/4}}$
$+\frac{L_{1}\log(\sqrt{n\log(d_{x})})}{\sqrt{T\log(d_{x})}}+\frac{L_{0}\sqrt{T\log(1/\delta)}\log^{3/2}(d_{x})}{n\varepsilon}.$
Hence letting $T=\min\left\{n,\frac{n\varepsilon}{\log(d_{x})\sqrt{\log(1/\delta)}}\right\}$ gives the claimed rate for second-order-smooth functions. Now, we look at the case where the partial derivates are not necessarily smooth. Now, we have
$\mathbb{E}[F_{\mathcal{D}}(w^{T})-F_{\mathcal{D}}(x)]$
$\leq\frac{\log(d_{x})}{\tau T}+\frac{\tau L_{0}^{2}}{2}+\frac{1}{T}\sum_{t\in[T]}\mathbb{E}\left[\langle\nabla F_{\mathcal{D}}(w^{t})-g^{t},x^{t}-x\rangle\right]$
$\lesssim\frac{\log(d_{x})}{\tau T}+\tau L_{0}^{2}+\frac{1}{T}\sum_{t\in[T]}\left(\frac{L_{1}}{\sqrt{K}}+\mathbbm{1}_{(t>q)}\left[\frac{4L_{1}q}{q\lfloor t/q\rfloor+1}+\frac{L_{1}^{2}q\tau}{\sqrt{\log(d_{x})}K^{3/2}}+\frac{(L_{0}^{2}+L_{1}^{2})\sqrt{\log(d_{x})}q\tau}{\sqrt{K}}\right]\right)$
$\lesssim\frac{\log(d_{x})}{\tau T}+\tau\Big{[}L_{0}^{2}+\frac{L_{1}^{2}q}{\sqrt{\log(d_{x})}K^{3/2}}+\frac{(L_{0}^{2}+L_{1}^{2})\sqrt{\log(d_{x})}q}{\sqrt{K}}\Big{]}+\frac{L_{1}}{\sqrt{K}}+\frac{1}{T}\sum_{t=q+1}^{T}\left[\frac{L_{1}q}{q\lfloor t/q\rfloor+1}\right]$
@@
By Corollary 3.1 the sparsification error is $\mathbb{E}[F_{\mathcal{D}}(\hat{w}^{T})-F_{\mathcal{D}}(w^{T})]\lesssim\frac{L_{1}}{K}$, which implies
$\mathbb{E}[F_{\mathcal{D}}(\hat{w}^{T})-F_{\mathcal{D}}(x)]\lesssim\frac{\log(d_{x})}{\tau T}+\tau\Big{[}L_{0}^{2}+\frac{L_{1}^{2}q}{\sqrt{\log(d_{x})}K^{3/2}}+\frac{(L_{0}^{2}+L_{1}^{2})\sqrt{\log(d_{x})}q}{\sqrt{K}}\Big{]}+\frac{L_{1}}{\sqrt{K}}+\frac{L_{1}q\log(T/q)}{T}.$
@@
$\tau\eqsim\min\left\{\sqrt{\frac{\log(d_{x})}{(L_{0}^{2}+(L_{0}^{2}+L_{1}^{2})q\sqrt{\log(d_{x})/K}+L_{1}^{2}q/[\sqrt{\log(d_{x})}K^{3/2}])T}},\frac{1}{L_{0}q},\frac{n\varepsilon}{TL_{0}\sqrt{(TK/q+qK)\log(1/\delta)}}\right\}$ leads to the convergence rate claimed for first-order-smooth functions. ∎
Appendix G Algorithm for Statistical Synthetic Data Generation
We introduced the Statistical Synthetic Data Generation problem in Appendix B. Recall its SSP formulation (3)
$\min_{x\in\Delta_{|\mathcal{Z}|}}\max_{y\in\Delta_{|\mathcal{Q}|}}\mathbb{E}_{z\sim\mathcal{D}}\Big{[}\sum_{j\in[|\mathcal{Q}|]}y_{j}(q_{j}(z)-\langle q_{j},x\rangle)\Big{]}.$
We denote the objective function by $f(x,y;z)=\sum_{j\in[|\mathcal{Q}|]}y_{j}(q_{j}(z)-\langle q_{j},x\rangle)$. This objective is convex-concave, since it is linear on $x$ and on $y$. It is easy to check that the it is also Lipschitz and smooth w.r.t $\|\cdot\|_{1}$. Furthemore, it is quadratic, so we could use Algorithm 1 directly. However, the algorithm below leverages the structure of the objective that we are optimizing in this specific formulation. The algorithm is the following
Algorithm 5 DP synthetic data generation with DP stochastic mirror descent and Maurey sparsification
0:  Dataset $S$, Number of steps $T$, step-sizes $\tau_{x},\tau_{y}$.
1:  $x^{1}=(1/|\mathcal{Z}|,...,1/|\mathcal{Z}|),y^{1}=(1/|\mathcal{Q}|,...,1/|\mathcal{Q}|)$
2:  for $t=1$ to $T$ do
3:     $\hat{y}^{t}\sim P_{y^{t}}$
4:     $g^{t}_{x}=\nabla_{x}f(x^{t},\hat{y}^{t};S)$
5:     $g^{t}_{y}=\nabla_{y}f(x^{t},\hat{y}^{t};S)$
6:     $x^{t+1}_{j}\propto x^{t}_{j}\exp\left(-\tau_{x}g^{t}_{x,j}\right),\quad\forall j\in[|\mathcal{Z}|]$
7:     $y^{t+1}_{j}\propto y^{t}_{j}\exp\left(\tau_{y}g^{t}_{y,j}\right),\quad\forall j\in[|\mathcal{Q}|]$
8:  end for
9:  $\bar{x}=\frac{1}{T}\sum_{t=1}^{T}x^{t}$
[…]
11:  return  $\tilde{x}$
We obtain the following theorem.
Theorem G.1.
Let $10<\varepsilon,\delta<$ 1. If $|\mathcal{Q}|\leq|\mathcal{Z}|^{C}$, Algorithm 5 with $T=\frac{6n\varepsilon}{16\sqrt{2\log(1/\delta)}\log(|{\cal Q}|)},\tau_{x}=\sqrt{\frac{\log(|\mathcal{Z}|)}{9T}},\tau_{y}=\frac{\log(|{\cal Q}|)}{6\sqrt{\log(|\mathcal{Z}|)T}}$ constructs an $(\varepsilon,\delta)$-DP synthetic dataset $\tilde{S}$ of $n$ samples, such that
$\mathbb{E}_{\mathcal{A},S}\Big{[}\max_{q\in\mathcal{Q}}|\mathbb{E}_{z\sim\mathcal{D}}[q(z)]-q(\tilde{S})|\Big{]}\leq 20\sqrt{\frac{\log(|{\cal Q}|)}{n}}+2(12+C/3)\frac{(\log(1/\delta)\log(|\mathcal{Z}|))^{1/4}\sqrt{\log(|{\cal Q}|)}}{\sqrt{n\varepsilon}}.$
Furthermore, this rate is optimal.
The proof of Theorem G.1 follows the same ideas we used for analyzing Algorithm 1.
Proof.
We prove the optimality, privacy and convergence rate claimed in the theorem seperately.
Optimality: The optimality of the rate is a folk result. The private lower bound follows from [Bun et al., 2014] combined with a resampling argument from [Bassily et al., 2019]. The non-private term comes from an application of Fano’s inequality [Cover and Thomas, 2006].
Privacy Analysis: Since $x^{1},y^{1}$ are trivially private, then $\tilde{y}^{1}$ is also private, as it is sampled according to the (private) probability vector $y^{1}$. Next, we have that $x^{2}$ is immediately private, as it only depends on $x_{1}$ and $\hat{y}_{1}$, and not on the data. However, $g_{y}^{1}$ depends on the data, so in order to privatize $y^{2}$ we create a private version of it using the the exponential mech as we did in Algorithm 1. Since we are using a full batch to estimate $g^{t}_{y}$, the sensitivity of the score function $s^{t}(S,j)=-\tau\left(\sum_{i=1}^{t}g^{i}_{y,j}\right)$ is $\frac{2TL_{0}\tau}{n}$. Now, we have that $x^{2}$ and $\hat{y}^{2}$ are both $\frac{4TL_{0}\tau}{n}$-DP. We inductively continue this argument, and using the advanced composition (Theorem A.1) to conclude that the average $\bar{x}$ is $(\varepsilon,\delta)$-DP so long as
$\frac{4TL_{0}\tau_{y}}{n}\leq\frac{\varepsilon}{2\sqrt{2T\log(1/\delta)}}.$
This condition holds under the parameter setting of Theorem G.1. The privacy of the output $\tilde{x}$ follows from post-processing of DP.
Accuracy Analysis: We start by noticing that
$\mathbb{E}\left[\max_{q\in{\cal Q}}|\mathbb{E}_{z}[q(z)]-q(\tilde{S})|\right]$ $=\mathbb{E}\left[\max_{y\in\Delta_{|{\cal Q}|}}\sum_{j\in[|{\cal Q}|]}y_{j}(\mathbb{E}_{z}[q(z)]-q(\tilde{S}))\right]$
$=\mathbb{E}\left[\max_{y\in\Delta_{|{\cal Q}|}}F_{\mathcal{D}}(\tilde{x},y)\right]$
$\leq\mathbb{E}\left[\max_{y\in\Delta_{|{\cal Q}|},x\in\Delta_{|\mathcal{Z}|}}(F_{\mathcal{D}}(\tilde{x},y)-F_{\mathcal{D}}(x,\bar{y})).\right]$
where the inequality comes from the fact that $F_{\mathcal{D}}(x_{\mathcal{D}},\bar{y})=0$ (where $x_{\mathcal{D}}$ is the probability vector associated with the unknown distribution $\mathcal{D}$ supported on $\mathcal{Z}$), so $\min_{x\in\Delta_{|\mathcal{Z}|}}F_{\mathcal{D}}(x,\bar{y})\leq 0$. Hence, it suffices that we bound $\mathbb{E}\left[\operatorname{Gap}(\tilde{x},\bar{y})\right]$. To this end, first we bound $\mathbb{E}\left[\operatorname{Gap}(\bar{x},\bar{y})\right]$. Analogously to the proof of Lemma 4.1, we note that
$T[F_{\mathcal{D}}(\bar{x},v)-F_{\mathcal{D}}(w,\bar{y})]\leq\operatorname{Regret}_{T}^{x}(w)+\operatorname{Regret}_{T}^{y}(v)+\sum_{t=1}^{T}\langle\Phi_{\mathcal{D}}(x^{t},y^{t})-g^{t},(x^{t},y^{t})-(w,v)\rangle$
$\leq\frac{\log(|\mathcal{Z}|)}{\tau_{x}}+\frac{\tau_{x}}{2}\sum_{t\in[T]}\|g^{t}_{x}\|_{\infty}^{2}+\frac{\log(|{\cal Q}|)}{\tau_{y}}+\frac{\tau_{y}}{2}\sum_{t\in[T]}\|g^{t}_{y}\|_{\infty}^{2}+\sum_{t=1}^{T}\langle\Phi_{\mathcal{D}}(x^{t},y^{t})-g^{t},(x^{t},y^{t})-(w,v)\rangle.$
Next, we have
$\sum_{t=1}^{T}\langle\Phi_{\mathcal{D}}(x^{t},y^{t})-g^{t},(x^{t},y^{t})-(w,v)\rangle=\sum_{t=1}^{T}\langle\nabla_{x}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{x},x^{t}-w\rangle+\langle-\nabla_{y}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{y},y^{t}-v\rangle$
$\leq\sum_{t=1}^{T}\langle\nabla_{x}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{x},x^{t}-w\rangle+2\|-\nabla_{y}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{y}\|_{\infty}$
$\leq\sum_{t=1}^{T}\langle(\nabla_{x}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{x},x^{t}-w^{t}\rangle)+\frac{\log(|\mathcal{Z}|)}{\tau_{x}}+\sum_{t=1}^{T}\tau_{x}\|\nabla_{x}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{x}\|_{\infty}^{2}+2\|\nabla_{y}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{y}\|_{\infty},$
where the first inequality follows from Holder’s inequality and the second from Lemma A.3, where $(v^{t})_{t\in[T]}$ are defined analogously to Lemma 4.1. Using these bounds above together with the fact that $L_{0}=2$, and maximizing over $w\in\Delta_{|\mathcal{Z}|},v\in\Delta_{|{\cal Q}|}$, we arrive to
$\operatorname{Gap}(\bar{x},\bar{y})$
$\leq\frac{2\log(|\mathcal{Z}|)}{\tau_{x}T}+\frac{\log(|{\cal Q}|)}{\tau_{y}T}+18\tau_{x}+2\tau_{y}+\frac{1}{T}\sum_{t=1}^{T}\langle\nabla_{x}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{x},x^{t}-w^{t}\rangle+2\|-\nabla_{y}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{y}\|_{\infty}.$
Taking expectations, we obtain
$\mathbb{E}[\operatorname{Gap}(\bar{x},\bar{y})]\leq\frac{2\log(|\mathcal{Z}|)}{\tau_{x}T}+\frac{\log(|{\cal Q}|)}{\tau_{y}T}+18\tau_{x}+2\tau_{y}+\frac{2}{T}\sum_{t=1}^{T}\mathbb{E}[\|-\nabla_{y}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{y}\|_{\infty}].$
We bound the standard deviation of $g^{t}_{y}$ using vector martingale concentration in $\ell_{\infty}$ norm (see Lemma A.2). More concretely, denoting $Q(z)=(q_{1}(z),...,q_{|{\cal Q}|}(z))$ and $Q(S)=\frac{1}{n}\sum_{j\in[n]}Q(z_{i_{j}})$,
$\mathbb{E}[\|-\nabla_{y}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{y}\|_{\infty}]$ $=\mathbb{E}[\|Q(S)-\mathbb{E}_{z}[Q(z)]\|_{\infty}]$
$\leq\sqrt{\mathbb{E}[\|Q(S)-\mathbb{E}_{z}[Q(z)]\|_{\infty}^{2}]}$
$=\sqrt{\mathbb{E}\left[\left\|\sum_{j\in[n]}\frac{Q(z_{i_{j}})-\mathbb{E}_{z}[Q(z)]}{n}\right\|_{\infty}^{2}\right]}$
$\leq\sqrt{2e\log(|{\cal Q}|)\sum_{j\in[n]}\left\|\frac{Q(z_{i_{j}})-\mathbb{E}_{z}[Q(z)]}{n}\right\|_{\infty}^{2}}$
$\leq 5\sqrt{\frac{\log(|{\cal Q}|)}{n}},$
where the first inequality follows from Jensen’s inequality and the second from Lemma A.2. We have proved the following bound for the expected gap
$\mathbb{E}[\operatorname{Gap}(\bar{x},\bar{y})]\leq\frac{2\log(|\mathcal{Z}|)}{\tau_{x}T}+\frac{\log(|{\cal Q}|)}{\tau_{y}T}+18\tau_{x}+2\tau_{y}+10\sqrt{\frac{\log(|{\cal Q}|)}{n}}.$
Setting $T=\frac{6n\varepsilon}{16\sqrt{2\log(1/\delta)}\log(|{\cal Q}|)},\tau_{x}=\sqrt{\frac{\log(|\mathcal{Z}|)}{9T}},\tau_{y}=\frac{\log(|{\cal Q}|)}{6\sqrt{\log(|\mathcal{Z}|)T}}$ and using $\log(|{\cal Q}|)\leq C\log(|\mathcal{Z}|)$, we obtain
$\mathbb{E}[\operatorname{Gap}(\bar{x},\bar{y})]\leq 10\sqrt{\frac{\log(|{\cal Q}|)}{n}}+2(12+C/3)\frac{(\log(1/\delta)\log(|\mathcal{Z}|))^{1/4}\sqrt{\log(|{\cal Q}|)}}{\sqrt{n\varepsilon}}.$
Also, following the proof of Lemma 4.2 together with the fact that $L_{1}=2$, we can show that
$\mathbb{E}[\operatorname{Gap}(\tilde{x},\bar{y})-\operatorname{Gap}(\bar{x},\bar{y})]\leq\frac{4}{n}+\frac{8}{\lambda n}+\lambda\log(|{\cal Q}|).$
Using $\lambda=\sqrt{8}/\sqrt{\log(|{\cal Q}|)T}$ gives $\mathbb{E}[\operatorname{Gap}(\tilde{x},\bar{y})-\operatorname{Gap}(\bar{x},\bar{y})]\leq\frac{4}{n}+6\sqrt{\frac{\log(|{\cal Q}|)}{n}}$. Finally, we conclude that
$\mathbb{E}\left[\max_{q\in{\cal Q}}|\mathbb{E}_{z}[q(z)]-q(\tilde{S})|\right]\leq\mathbb{E}[\operatorname{Gap}(\tilde{x},\bar{y})]\leq 20\sqrt{\frac{\log(|{\cal Q}|)}{n}}+2(12+C/3)\frac{(\log(1/\delta)\log(|\mathcal{Z}|))^{1/4}\sqrt{\log(|{\cal Q}|)}}{\sqrt{n\varepsilon}}.$

### v2

@@ 4.2 Convergence Analysis of Algorithm 5
$\,+2\tau q\|\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t})\|_{\infty}^{2}+2\tau\sum_{k=t_{0}}^{t-1}\mathbb{E}\left[\|\nabla f(\hat{w}^{t_{0}};B^{k})-\nabla f(w^{t_{0}};B^{k})\|_{\infty}^{2}\right]$
$\lesssim\begin{cases}\frac{L_{1}}{\sqrt{K}}+q\tau\big[\frac{L_{1}^{2}}{\sqrt{\log(d_{x})}K^{3/2}}+\frac{(L_{0}^{2}+L_{1}^{2})\sqrt{\log(d_{x})}}{\sqrt{K}}\big]&\text{for first-order-smooth $f(\cdot;z)$}\\ \frac{L_{2}}{K}+q\tau\big[\frac{L_{2}^{2}}{K^{2}}+\frac{L_{1}^{2}\log(d_{x})}{K}\big]&\text{for second-order-smooth $f(\cdot;z)$.}\end{cases}$
And recall that for $t\leq q$,
$\mathbb{E}[\langle\nabla F_{\mathcal{D}}(w^{t})-\nabla f(\hat{w}^{t};B^{t}),x^{t}-x\rangle]\leq\begin{cases}\frac{8L_{1}}{\sqrt{K}}&\text{for first-order-smooth $f(\cdot;z)$}\\ \frac{4L_{2}}{K}&\text{for second-order-smooth $f(\cdot;z)$.}\end{cases}$
Finally, we plug this bounds into our excess risk bound. We start considering the case of second-order-smooth functions. We bound $\mathbb{E}[F_{\mathcal{D}}(w^{T})-F_{\mathcal{D}}(x)]$ by
$\frac{\log(d_{x})}{\tau T}+\frac{\tau L_{0}^{2}}{2}+\frac{1}{T}\sum_{t\in[T]}\mathbb{E}\left[\langle\nabla F_{\mathcal{D}}(w^{t})-g^{t},x^{t}-x\rangle\right]$
$\textstyle\lesssim\frac{\log(d_{x})}{\tau T}+\tau L_{0}^{2}+\frac{1}{T}\sum_{t\in[T]}\Big(\frac{L_{2}}{K}+\mathbbm{1}_{(t>q)}\Big[\frac{4L_{1}q}{q\lfloor t/q\rfloor+1}+q\frac{L_{2}^{2}\tau}{K^{2}}+q\frac{L_{1}^{2}\log(d_{x})\tau}{K}\Big]\Big)$
$\textstyle\lesssim\frac{\log(d_{x})}{\tau T}+\tau\Big[L_{0}^{2}+\frac{L_{1}^{2}q\log(d_{x})}{K}+\frac{L_{2}^{2}q}{K^{2}}\big]+\frac{L_{2}}{K}+\frac{1}{T}\sum_{t=q+1}^{T}\frac{L_{1}q}{q\lfloor t/q\rfloor+1}$
(11) $\lesssim\frac{\log(d_{x})}{\tau T}+\tau\Big[L_{0}^{2}+\frac{L_{1}^{2}q\log(d_{x})}{K}+\frac{L_{2}^{2}q}{K^{2}}\big]+\frac{L_{2}}{K}+\frac{L_{1}q\log(T/q)}{T},$
where the last line follows from $\sum_{t=q+1}^{T}\left[\frac{1}{q\lfloor t/q\rfloor+1}\right]\leq q\left(\frac{1}{q}+\frac{1}{2q}+...+\frac{1}{(T/q)q}\right)\lesssim\log\big(\frac{T}{q}\big)$. Now, by Theorem 4.3, and recalling that the batch size is $B=n/T$, we have a constraint on $\tau>0$ in order to satisfy privacy. We also have the constraint of $\psi\leq 1$, which holds true only when $\tau\leq 1/(8qL_{0})$. Optimizing (11) under these constraints, we have that we should choose $\tau$ of order
$\min\bigg\{$ $\sqrt{\frac{\log(d_{x})}{(L_{0}^{2}+L_{1}^{2}q\log(d_{x})/K+L_{2}^{2}q/K^{2})T}},\frac{1}{L_{0}q},\frac{n\varepsilon}{TL_{0}\sqrt{(TK/q+qK)\log(1/\delta)}}\bigg\}.$
With this choice we obtain
@@
$\mathbb{E}[F_{\mathcal{D}}(\hat{w}^{T})-F_{\mathcal{D}}(x)]\lesssim(L_{0}+L_{1})\sqrt{\frac{\log(d_{x})}{T}}\\ \textstyle+\frac{L_{2}}{T^{3/4}\log(d_{x})^{1/4}}+\frac{L_{1}+L_{2}}{\sqrt{T\log(d_{x})}}+\frac{L_{1}\log(\sqrt{n\log(d_{x})})}{\sqrt{T\log(d_{x})}}+\frac{L_{0}\sqrt{T\log(1/\delta)}\log^{3/2}(d_{x})}{n\varepsilon}\\ \textstyle\lesssim(L_{0}+L_{1})\sqrt{\frac{\log(d_{x})}{T}}+\frac{L_{2}}{\sqrt{T}\log(d_{x})^{1/4}}+\frac{L_{1}\log(\sqrt{n\log(d_{x})})}{\sqrt{T\log(d_{x})}}+\frac{L_{0}\sqrt{T\log(1/\delta)}\log^{3/2}(d_{x})}{n\varepsilon}.$
Hence letting $T=\min\Big\{n,\frac{n\varepsilon}{\log(d_{x})\sqrt{\log(1/\delta)}}\Big\}$ gives the claimed rate for second-order-smooth functions. Now, we look at the case where the partial derivates are not necessarily smooth. Now, we can bound $\mathbb{E}[F_{\mathcal{D}}(w^{T})-F_{\mathcal{D}}(x)]$ by
$\frac{\log(d_{x})}{\tau T}+\frac{\tau L_{0}^{2}}{2}+\frac{1}{T}\sum_{t\in[T]}\mathbb{E}\left[\langle\nabla F_{\mathcal{D}}(w^{t})-g^{t},x^{t}-x\rangle\right]$
$\lesssim\frac{\log(d_{x})}{\tau T}+\tau L_{0}^{2}+\frac{1}{T}\sum_{t\in[T]}\bigg(\frac{L_{1}}{\sqrt{K}}+\mathbbm{1}_{(t>q)}\bigg[\frac{4L_{1}q}{q\lfloor t/q\rfloor+1}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\frac{L_{1}^{2}q\tau}{\sqrt{\log(d_{x})}K^{3/2}}+\frac{(L_{0}^{2}+L_{1}^{2})\sqrt{\log(d_{x})}q\tau}{\sqrt{K}}\bigg]\bigg)$
$\lesssim\frac{\log(d_{x})}{\tau T}+\tau\Big[L_{0}^{2}+\frac{L_{1}^{2}q}{\sqrt{\log(d_{x})}K^{3/2}}+\frac{(L_{0}^{2}+L_{1}^{2})\sqrt{\log(d_{x})}q}{\sqrt{K}}\Big]$
$\qquad\qquad+\frac{L_{1}}{\sqrt{K}}+\frac{1}{T}\sum_{t=q+1}^{T}\left[\frac{L_{1}q}{q\lfloor t/q\rfloor+1}\right]$
@@
By Corollary 2.3, the sparsification error is $\mathbb{E}[F_{\mathcal{D}}(\hat{w}^{T})-F_{\mathcal{D}}(w^{T})]\lesssim\frac{L_{1}}{K}$, which implies
$\mathbb{E}[F_{\mathcal{D}}(\hat{w}^{T})-F_{\mathcal{D}}(x)]\lesssim$ $\frac{\log(d_{x})}{\tau T}+\tau\Big[L_{0}^{2}+\frac{L_{1}^{2}q}{\sqrt{\log(d_{x})}K^{3/2}}+\frac{(L_{0}^{2}+L_{1}^{2})\sqrt{\log(d_{x})}q}{\sqrt{K}}\Big]$
$\quad+\frac{L_{1}}{\sqrt{K}}+\frac{L_{1}q\log(T/q)}{T}.$
@@
$\tau\!\eqsim\!\!\min\!\left\{\sqrt{\frac{\log(d_{x})}{(L_{0}^{2}+(L_{0}^{2}+L_{1}^{2})q\sqrt{\log(d_{x})/K}+L_{1}^{2}q/[\sqrt{\log(d_{x})}K^{3/2}])T}},\frac{1}{L_{0}q},\frac{n\varepsilon}{TL_{0}\sqrt{(TK/q+qK)\log(1/\delta)}}\right\}$ leads to the convergence rate claimed for first-order-smooth functions.
5 Conclusion and Future Work
We presented mirror descent based algorithms for DP-SSP in the polyhedral setting. Our algorithms are the first achieving polylogarithmic-in-the-dimension error beyond bilinear objectives. In fact, for second-order-smooth objectives our rates are optimal up to logarithmic factors. An important question remains: Are the optimal rates achievable in expectation for first-order-smooth functions? We leave this question for future work. We also apply our ideas to DP-SCO in the polyhedral setting, where we also leave open the problem of achieving optimal rates with mirror descent based algorithms.
6 Acknowledgments
T. González’s research was partially done as a Research Intern at Google DeepMind. C. Guzmán’s research was partially done as a Visiting Researcher at Google. C. Guzmán’s research was partially supported by INRIA Associate Teams project, ANID FONDECYT 1210362 grant, ANID Anillo ACT210005 grant, and National Center for Artificial Intelligence CENIA FB210017, Basal ANID. C. Paquette is a CIFAR AI chair, (MILA) and C. Paquette was supported by a NSERC Discovery Grant, NSERC CREATE grant, an external grant from Google, and FRQNT New University Researcher’s Start-Up Program.